What is the slope intercept form equation of this?
(1, -1), m= -3/4

Found 2 solutions by josgarithmetic, lwsshak3:
Answer by josgarithmetic(39617)

The form you want is y=mx+b.
You are given m, and a point on the line. You want to find the value of b. Solve the formula for b and use the given point to compute b.

Answer by lwsshak3(11628)

What is the slope intercept form equation of this?
(1, -1), m= -3/4
Slope intercept form of equation for a line: y=mx+b, m=slope, b=y-intercept
For given problem:
y=-(3/4)x+b
solve for b using given coordinates:
-1=-3*1/4+b
b=-1/4
equation of given line=-3x/4-1/4